# DEEP LEARNING-DRIVEN MALWARE CLASSIFICATION WITH API CALL SEQUENCE ANALYSIS AND CONCEPT DRIFT HANDLING

### v1

@@ Abstract
Keywords API calls  $\cdot$ Malware Classifier  $\cdot$ $n$-grams  $\cdot$ Portable Executable  $\cdot$ Fitness Score  $\cdot$ Mutation
@@ 1 Introduction
In the field of cybersecurity, the examination of Portable Executable (PE) files plays a crucial role in detecting and analyzing malware. This study introduces a novel method for dynamic PE malware analysis, leveraging advanced deep-learning techniques to enhance classification accuracy and efficiency. Our proposed model employs neural networks to categorize malware samples into distinct families based on their behaviour and intrinsic characteristics.
A standout feature of our approach is the integration of Genetic Algorithms to manage concept drift, which is a common challenge in malware detection due to the evolving nature of malicious software. Concept drift refers to the changes in the statistical properties of the target variable, necessitating continuous adaptation of the model to maintain its performance over time. By incorporating Genetic Algorithms, our classification system not only remains effective but also adapts dynamically to new and emerging threats, ensuring robust and reliable malware detection. This methodology underscores the importance of adaptability and resilience in cybersecurity strategies, paving the way for more advanced and responsive defense mechanisms against sophisticated cyber threats.
This paper aims at exploring the application of deep learning techniques with a combination of genetic algorithm approaches for dynamic malware analysis, focusing on $n$-gram API calls. We aim at leveraging deep learning techniques to identify patterns and behaviours in malware samples, enhancing our understanding and classification capabilities with concept drift.
The remaining sections of the chapter are structured as follows: Section 2 explains the basic concepts of concept drift, its types, and the role of genetic algorithms in concept drift handling and sandbox environments. Section 3 discusses the related work on concept drift. Section 4 details the data collection process. Section 5 presents our DL and genetic algorithms-based framework for malware classification, covering preprocessing, NLP, feature selection, malware classification phases, and concept drift handling. Section 6 describes our experimental setup, including obtaining an API key, data extraction, malware classification, and concept drift handling, and provides a detailed analysis of results obtained by using different deep learning algorithms combined with genetic algorithms. Section 7 compares our approach with state-of-the-art techniques in malware detection. Section 8 covers the threads to validity. Finally, Section 9 presents the conclusion and the future work of our paper.
2 Basic Concepts
In this section, we introduce the fundamental concepts and terminologies necessary to understand this paper, focusing on concept drift, genetic algorithms, and sandbox environments.
2.1 Concept-drift
Concept drift [1][2] refers to the phenomenon in which the statistical properties of a target variable, which a model is trying to predict, change over time in unforeseen ways. This can happen for various reasons, such as changes in the environment, the data source, or the relationships between variables. Concept drift can be categorised into several types based on the nature of the change:
@@ 2.1 Concept-drift
    Sudden Concept Drift: This type of drift occurs when there is an abrupt and significant change in the target variable’s distribution. It can be caused by events such as a sudden change in user behaviour or a change in the underlying process generating the data.
@@
    Incremental Concept Drift: Incremental drift refers to a gradual, continuous change in the target variable’s distribution over time. This type of drift can be challenging to detect, as the changes are subtle and accumulate over time.
@@
    Recurring Concept Drift: Recurring drift occurs when the target variable’s distribution follows a recurring pattern or cycle. For example, seasonal changes in user behaviour or periodic changes in the underlying process can lead to recurring drift [3].
@@
    Concept Drift by Context: This type of drift occurs when the relationship between the input variables and the target variable changes based on the context or conditions under which the data is generated. Contextual drift can be particularly challenging to model, as it requires capturing the contextual information along with the data.
@@
    Covariate Shift: Covariate shift refers to a change in the distribution of the input variables while the relationship between the input and target variables remains the same. This type of drift can be addressed by reweighting the data to match the new distribution of the input variables [4].
Detecting and adapting to concept drift is essential for maintaining the performance of machine learning models in dynamic environments. Various techniques, such as online learning, ensemble methods, and monitoring performance metrics over time, can be used to mitigate the impact of concept drift.
2.2 Techniques to Handle Concept Drift
Handling concept drift is crucial for maintaining the performance of machine learning models in dynamic environments. Below, we present some several techniques commonly used to address concept drift.
@@ 2.2 Techniques to Handle Concept Drift
    Re-training: Periodically retraining the model using the most recent data can help the model adapt to concept drift. This approach is effective for both sudden and incremental drift.
@@
    Ensemble Methods: Ensemble methods, such as bagging and boosting, can improve model robustness to concept drift by combining the predictions of multiple models trained on different subsets of data or with different algorithms.
@@
    Online Learning: Online learning techniques allow models to be updated continuously as new data becomes available. This approach is well-suited for handling incremental drift.
@@
    Change Detection Algorithms: Change detection algorithms, such as Cumulative Sum (CUSUM) and Page-Hinkley, can be used to detect concept drift by monitoring the model’s performance metrics over time.
@@
    Instance Weighting: Instance weighting techniques, such as Importance Weighted Cross Validation (IWCV), can be used to give more weight to recent data instances, helping the model adapt to incremental drift.
@@
    Feature Selection and Extraction: Adapting the feature set based on the changing data distribution can help improve the model’s performance in the presence of concept drift.
@@
    Windowing: Using a sliding window approach to train the model on the most recent data can help the model adapt to gradual changes in data distribution.
@@
    Dynamic Feature Adaptation: Modifying the feature representation based on the current data distribution can help the model better capture the underlying patterns in the data.
@@
    Memory-based Methods: Memory-based methods, such as k-nearest neighbours (KNN), can be adapted to store historical data and update the model as new data becomes available, allowing the model to adapt to concept drift.
Each of these techniques has its strengths and weaknesses, and the choice of technique depends on the specific characteristics of the data and the problem at hand.
2.3 Genetic Algorithms for Concept Drift
Genetic algorithms (GAs) [5] can be used to handle concept drift in machine-learning models. Genetic algorithms are optimisation algorithms inspired by the process of natural selection and genetics. They can be applied to adapt the model to changing environments, including concept drift scenarios. Here’s how GAs can be used for concept drift:
@@ 2.3 Genetic Algorithms for Concept Drift
    Feature Selection: Genetic algorithms can be used to dynamically select the most relevant features for the current concept. This can help the model adapt to changing data distributions.
@@
    Model Selection: GAs can be used to evolve the structure of the model, such as the type of model or the hyperparameters, to better fit the current concept.
@@
    Ensemble Generation: GAs can be used to create ensembles of models with different structures or parameters, allowing the ensemble to adapt to changing concepts.
@@
    Hyperparameter Optimization: GAs can be used to optimize the hyperparameters of the model to improve its performance in the current concept.
@@
    Instance Selection: GAs can be used to select a subset of instances from the historical data that are most relevant to the current concept, helping the model adapt more effectively.
By using genetic algorithms in these ways, machine learning models can be made more robust to concept drift, allowing them to maintain high performance even as the data distribution changes over time.
2.4 Sandbox
A sandbox is a security mechanism used in software development, testing, and cybersecurity to isolate running programs or processes from the rest of the system. Inspired by children’s sandbox play areas, it provides a controlled environment where untrusted or potentially harmful code can be executed without risking the underlying system [6].
2.5 Working Principle of a Sandbox
  1. 1.
    Resource Control: Sandboxes limit the resources available to processes, such as files, network resources, memory, and CPU usage, to mitigate potential damage from malicious code.
  2. 2.
    Monitoring and Analysis: Sandboxes include monitoring capabilities to log and analyze system calls, network activity, file operations, and other interactions to detect suspicious behavior.
  3. 3.
    Dynamic Analysis: Sandboxes use real-time dynamic analysis to detect and respond to threats, identifying attempts to exploit vulnerabilities or execute malicious actions [7].
  4. 4.
    Containment: Upon detecting malicious behavior, sandboxes can terminate the offending process or revert changes to the system to contain the threat.
2.6 Cuckoo Sandbox
Cuckoo Sandbox [^1] https://github.com/cuckoosandbox/ [8] [9] is an open-source automated malware analysis system designed to analyse suspicious files and URLs within a controlled environment. It utilises a virtualised environment to execute potentially malicious code safely and provides detailed analysis reports for further investigation.
2.6.1 System Requirements
Before installing Cuckoo Sandbox, we have to ensure that our system meets the following requirements:
  • Availability of Linux-based operating system (e.g., Ubuntu, Debian, CentOS)
  • Adequate hardware resources (CPU, RAM, Disk space) for running virtual machines
  • Availability of Virtualization software (e.g., VirtualBox, VMware, KVM)
  • Python 3.x installed

### v2

@@ Abstract
Keywords:
API calls Malware Classifier $n$-grams Portable Executable Fitness Score Mutation
journal: Journal Information
@@ 1 Introduction
In the ever-evolving landscape of cybersecurity, the analysis of Portable Executable (PE) files—a prevalent format for executable programs in Windows operating systems—stands as a cornerstone for detecting and mitigating malware threats. PE files encapsulate critical structural and behavioral information, making them a prime target for malicious exploitation by cybercriminals. As malware authors continuously refine their techniques to obfuscate code, bypass signature-based detection, and exploit system vulnerabilities, traditional static analysis methods falter, unable to keep pace with the sophistication and dynamism of these threats. This necessitates a shift toward dynamic analysis approaches that can adapt to the mutable nature of malware, a challenge compounded by the phenomenon of concept drift tsymbal2004problem .
Concept drift, defined as the temporal alteration in the statistical properties of a target variable, poses a significant hurdle in maintaining the efficacy of machine learning models for malware detection. In the context of cybersecurity, this drift manifests as evolving malware behaviors—such as changes in API call patterns, payload delivery mechanisms, or encryption strategies—driven by adversaries’ relentless innovation. Without adaptive mechanisms, pre-trained models quickly degrade, failing to recognize new variants or emerging families of malicious software lu2018learning . Motivated by this pressing need for adaptability, our study introduces a novel framework that synergistically combines advanced deep learning (DL) techniques with genetic algorithms (GAs) to enhance the accuracy, efficiency, and resilience of dynamic PE malware analysis.
Our proposed model leverages the power of neural networks—specifically Artificial Neural Networks (ANNs), Convolutional Neural Networks (CNNs), and Recurrent Neural Networks (RNNs)—to classify malware samples into distinct families based on their behavioral signatures, extracted primarily through $n$-gram analysis of API call sequences. These sequences, derived from dynamic execution in sandbox environments, encapsulate the runtime interactions of PE files with the operating system, offering a rich feature set for discerning malicious intent. However, the true innovation of our approach lies in its integration of GAs to address concept drift. Inspired by natural selection, GAs enable continuous model evolution by applying mutation operations and fitness-based selection to adapt feature sets and optimize classification performance in real time vivekanandan2011mining . This dual strategy ensures that our system not only excels in categorizing known malware but also dynamically adjusts to emerging threats, a capability critical for safeguarding systems in today’s volatile cyber landscape.
The motivation behind this research stems from the escalating arms race between malware developers and cybersecurity defenders. Static models, while computationally efficient, lack the flexibility to counter the rapid evolution of malware, as evidenced by studies highlighting performance degradation over time chen2023overkill . Conversely, purely DL-based approaches, though powerful in pattern recognition, often struggle with generalization across drifting data distributions without retraining jameel2020critical . By marrying DL’s robust feature extraction with GAs’ evolutionary adaptability, we aim to bridge this gap, offering a proactive, scalable solution that anticipates and responds to shifts in malware characteristics. This work builds on prior efforts to handle concept drift garcia2023effectiveness ; fernando2024fesad while pushing the boundaries of malware classification through a focus on API call sequences—a dynamic, behavior-driven indicator of malicious activity.
Our objectives are twofold: first, to harness deep learning techniques to uncover intricate patterns and behaviors within malware samples, thereby improving classification precision; and second, to enhance adaptability to concept drift, ensuring long-term effectiveness against sophisticated cyber threats. To this end, the paper explores the application of DL and GA methodologies for dynamic malware analysis, with a particular emphasis on $n$-gram API call analysis as a lens into malware behavior.
The remainder of this paper is structured as follows: Section 2 elucidates the foundational concepts of concept drift, its variants, and the pivotal role of GAs and sandbox environments in addressing it. Section 3 reviews related work, situating our contribution within the broader research landscape. Section 4 delineates the data collection and preprocessing pipeline critical to our experiments. Section 5 presents our integrated DL and GA-based framework, detailing its phases from preprocessing to concept drift handling. Section 6 describes the experimental setup—including API key acquisition, data extraction, and model training—and provides an in-depth analysis of results across diverse DL algorithms paired with GAs. Section 7 benchmarks our approach against state-of-the-art techniques, highlighting its strengths and novelty. Section 8 discusses threats to validity, ensuring a rigorous evaluation of our findings. Finally, Section 9 summarizes our contributions and outlines avenues for future research, reinforcing the transformative potential of our methodology in fortifying cybersecurity defenses.
2 Foundational Concepts for Malware Analysis
This section lays the groundwork for understanding the core methodologies employed in our study by elucidating the essential concepts and terminologies central to dynamic malware analysis. We focus on three intertwined pillars: concept drift, which captures the evolving nature of malware data; techniques to handle concept drift, which ensure model robustness; and genetic algorithms, which provide an adaptive framework for addressing these shifts. These concepts are critical to our proposed deep learning (DL) and genetic algorithm (GA)-based approach for classifying Portable Executable (PE) malware, particularly in the context of API call sequence analysis.
Concept Drift: Definition and Taxonomy
Concept drift refers to the phenomenon wherein the statistical properties of a target variable—the entity a predictive model aims to classify or forecast—undergo unforeseen changes over time tsymbal2004problem ; lu2018learning . In machine learning, this drift manifests as a discrepancy between the training data distribution $P(X,y)$ and the test-time distribution $P^{\prime}(X,y)$, where $X$ represents input features (e.g., API call sequences) and $y$ denotes the target labels (e.g., malware families). Formally, concept drift occurs when:
$P_{t}(X,y)\neq P_{t+1}(X,y),$
where $t$ and $t+1$ denote successive time steps. In the cybersecurity domain, such shifts are driven by factors like adversarial innovation (e.g., new obfuscation techniques), environmental changes (e.g., operating system updates), or data source variations (e.g., emerging malware strains). For malware detection, concept drift complicates the maintenance of model accuracy, as static models trained on historical data fail to generalize to novel threats.
To characterize this phenomenon, concept drift is categorized into distinct types based on the nature and tempo of the change:
@@ 2 Foundational Concepts for Malware Analysis
    Sudden Concept Drift: This occurs when the target distribution shifts abruptly, often triggered by a discrete event. For instance, the release of a new malware variant (e.g., WannaCry ransomware in 2017) could abruptly alter the distribution of malicious behaviors, rendering existing models obsolete overnight. Mathematically, this is a step change: $P_{t}(y|X)\to P_{t+1}(y|X)$ where $P_{t+1}$ differs significantly within a short timeframe.
@@
    Incremental Concept Drift: This type involves a gradual, continuous evolution of the target distribution. An example is the slow adaptation of trojan malware to employ more sophisticated encryption over months, making detection incrementally harder. This can be modeled as a smooth transition: $P_{t}(y|X)\to P_{t+1}(y|X)$ with small, cumulative changes.
@@
    Recurring Concept Drift: Here, the distribution exhibits cyclic or periodic patterns. Seasonal ransomware campaigns, such as those peaking during tax seasons, exemplify this, where $P_{t}(y|X)$ reverts to a prior state after a cycle (e.g., $P_{t+n}(y|X)\approx P_{t}(y|X)$) suarez2023survey . This requires models to recognize and adapt to temporal recurrence.
@@
    Concept Drift by Context: This arises when the relationship between inputs $X$ and outputs $y$ varies with contextual factors. For malware, context might include the target operating system version; a backdoor’s behavior may differ between Windows 7 and Windows 10, necessitating context-aware modeling.
@@
    Covariate Shift: Unlike other types, covariate shift involves a change in the input distribution $P(X)$ while the conditional distribution $P(y|X)$ remains stable nair2019covariate . In malware analysis, this might occur when API call frequencies shift due to software updates, yet the underlying malicious intent remains consistent. Mitigation often involves reweighting samples via importance sampling: $w(X)=\frac{P_{\text{new}}(X)}{P_{\text{old}}(X)}$.
Detecting and adapting to concept drift is paramount in dynamic environments like malware detection, where failure to adapt leads to degraded performance (e.g., increased false negatives). Techniques range from passive monitoring of prediction errors to proactive model updates, as detailed next.
Techniques to Handle Concept Drift
Maintaining model efficacy amidst concept drift requires a suite of adaptive strategies tailored to the drift’s characteristics. Below, we outline key techniques, emphasizing their applicability to malware classification:
@@
    Re-training: Periodic retraining updates the model with fresh data, resetting parameters to reflect $P_{t+1}(X,y)$. For sudden drift (e.g., a new worm outbreak), full retraining on recent samples restores accuracy, while incremental drift benefits from partial updates. However, this approach demands computational resources and timely data availability.
@@
    Ensemble Methods: Techniques like bagging and boosting combine multiple models (e.g., decision trees, neural networks) trained on diverse data subsets or time windows. For malware, an ensemble might include classifiers for old and new variants, weighted by recency or performance, enhancing robustness to recurring drift kuncheva2004classifier.
@@
    Online Learning: This enables continuous parameter updates as new data arrives, ideal for incremental drift. Algorithms like stochastic gradient descent (SGD) adjust weights incrementally: $w_{t+1}=w_{t}-\eta\nabla L(x_{t},y_{t})$, where $\eta$ is the learning rate and $L$ is the loss. In cybersecurity, this suits streaming malware data from live systems.
@@
    Change Detection Algorithms: Methods such as Cumulative Sum (CUSUM) or Page-Hinkley test monitor performance metrics (e.g., error rate) for drift signals. CUSUM tracks cumulative deviations: $S_{t}=\max(0,S_{t-1}+(x_{t}-\mu))$, triggering retraining when $S_{t}$ exceeds a threshold. This is vital for sudden drift detection in malware campaigns.
@@
    Instance Weighting: Techniques like Importance Weighted Cross-Validation (IWCV) assign higher weights to recent instances, adapting to incremental drift. For a sample $(x_{i},y_{i})$, the weight $w_{i}=e^{-\lambda(t-t_{i})}$ prioritizes newer data, where $\lambda$ controls decay rate.
@@
    Feature Selection and Extraction: Dynamically updating features (e.g., selecting new API calls) aligns the model with shifting distributions, critical for covariate shift in malware datasets.
@@
    Windowing: A sliding window trains on the most recent $n$ samples, discarding older data. For gradual drift, a fixed-size window (e.g., 1000 samples) ensures relevance, though it risks losing long-term patterns.
@@
    Dynamic Feature Adaptation: This modifies feature representations (e.g., $n$-gram weights) based on current data, enhancing pattern capture amidst drift.
@@
    Memory-Based Methods: Approaches like k-Nearest Neighbors (k-NN) store historical data, querying it for predictions. Adaptive k-NN adjusts the neighbor set dynamically, balancing memory and recency for recurring drift.
Each strategy offers trade-offs: online learning excels in real-time settings but may overfit noise, while ensembles provide stability at higher computational cost. The choice depends on drift type, data volume, and application constraints, with our study favoring GAs for their evolutionary adaptability.
Genetic Algorithms for Concept Drift
Genetic algorithms (GAs), inspired by Darwinian evolution, are optimization heuristics that emulate natural selection through mutation, crossover, and selection vivekanandan2011mining . In machine learning, GAs address concept drift by evolving solutions to align with changing environments. Operating on a population of candidate solutions (e.g., feature sets, models), GAs iterate as follows: 1. Initialization: Generate a random population. 2. Fitness Evaluation: Score individuals using a fitness function (e.g., classification accuracy). 3. Selection: Favor high-fitness individuals (e.g., via tournament selection). 4. Crossover: Combine parent solutions to produce offspring. 5. Mutation: Introduce random changes to maintain diversity. 6. Replacement: Update the population with new generations.
For concept drift, GAs adapt models dynamically:
@@
    Feature Selection: GAs identify relevant features (e.g., API $n$-grams) for the current drift state, evolving a chromosome of binary feature masks (1 = selected, 0 = discarded). Fitness might be F1-score on recent data.
@@
    Model Selection: GAs evolve model architectures (e.g., neural network layers) or types (e.g., CNN vs. RNN), optimizing for current $P(y|X)$.
@@
    Ensemble Generation: GAs create diverse model ensembles, each tailored to a drift phase, combining predictions via weighted voting.
@@
    Hyperparameter Optimization: GAs tune parameters (e.g., learning rate $\eta$, dropout rate) to maximize performance, encoding them as real-valued genes.
@@
    Instance Selection: GAs select representative historical instances, reducing noise and aligning training data with $P_{t+1}(X,y)$.
In malware analysis, GAs excel by mutating API call features to mimic evolving threats (e.g., new system call patterns), ensuring robustness. For example, a fitness function $F=\text{Accuracy}+\alpha\cdot\text{Stability}$ (where $\alpha$ balances performance and drift resilience) guides evolution. This adaptability distinguishes GAs from static retraining, making them integral to our framework, as detailed in Section 5.
2.1 Sandboxing in Malware Analysis
In the domain of cybersecurity and software engineering, a sandbox serves as a pivotal security mechanism designed to isolate and execute untrusted or potentially malicious programs in a controlled environment, safeguarding the underlying system from harm. Drawing an analogy from a child’s sandbox—where activities are confined to a safe, bounded space—this technique creates a virtualized or restricted execution context that prevents rogue processes from interacting directly with critical system resources vasilescu2014practical . Sandboxes are widely employed in software development, testing, and, most notably, malware analysis, where they enable the safe detonation and observation of suspicious code. This isolation is particularly crucial for analyzing Portable Executable (PE) files, which are frequent vectors for malware due to their executable nature in Windows environments.
The operational foundation of a sandbox rests on a multi-faceted approach to resource management, monitoring, and threat containment, ensuring both security and analytical depth. First, resource control is enforced by imposing strict limits on the resources available to the sandboxed process. These restrictions encompass file system access (e.g., read/write permissions confined to a virtual directory), network connectivity (e.g., blocking outbound traffic or rerouting it to a honeypot), memory allocation (e.g., capping usage to prevent buffer overflows), and CPU utilization (e.g., throttling to mitigate denial-of-service risks). By constraining these resources, the sandbox minimizes the potential damage that malicious code could inflict, even if it attempts to exploit vulnerabilities afianian2019malware .
Second, monitoring and analysis form the backbone of a sandbox’s diagnostic capabilities. Advanced sandboxes are equipped with instrumentation to log and scrutinize a wide array of system interactions, including system calls (e.g., NtCreateFile, NtWriteFile), network activity (e.g., DNS queries, TCP/UDP packet flows), file operations (e.g., creation, modification, deletion), and registry manipulations (e.g., key creation or value changes). These logs are analyzed in real time or post-execution to identify behavioral signatures indicative of malice, such as attempts to escalate privileges, inject code into other processes, or establish command-and-control (C2) communications. This dynamic analysis contrasts with static analysis by capturing runtime behavior, offering a more comprehensive view of a malware sample’s intent and capabilities jamalpur2018dynamic .
Third, the sandbox employs dynamic analysis to detect and respond to threats as they unfold. Unlike static methods that disassemble code without execution, dynamic analysis executes the sample within the sandbox, observing its interactions with the emulated system. Techniques such as API hooking, kernel-level monitoring, and behavioral heuristics enable the identification of exploits (e.g., zero-day vulnerabilities) or malicious actions (e.g., ransomware encryption routines). This real-time scrutiny is computationally intensive but invaluable for uncovering obfuscated or polymorphic malware that evades static detection afianian2019malware .
Finally, containment mechanisms ensure that detected threats are neutralized effectively. Upon identifying suspicious behavior—such as a process attempting to overwrite the Master Boot Record (MBR) or connect to a known malicious IP—the sandbox can terminate the process, roll back system changes using snapshot restoration (common in virtualized sandboxes), or quarantine the sample for further study. This containment not only prevents harm but also preserves the integrity of the analysis environment, allowing repeated executions under varied conditions.
A prominent implementation of these principles is Cuckoo Sandbox, an open-source, automated malware analysis system accessible at https://github.com/cuckoosandbox/ cuckoo_repo . Cuckoo Sandbox exemplifies the sandbox paradigm by providing a robust framework for analyzing suspicious files and URLs within a virtualized setting. It leverages virtualization technologies (e.g., VirtualBox, VMware, or KVM) to instantiate isolated guest environments—typically Windows instances—where PE files are executed safely. Cuckoo’s architecture comprises a host system that orchestrates analysis and a guest virtual machine (VM) where the malware is detonated. During execution, it captures detailed telemetry, including API call sequences, network traffic (via packet capture), and system modifications, which are compiled into comprehensive JSON reports for subsequent investigation jamalpur2018dynamic . These reports are instrumental in our study, as they provide the $n$-gram API call data driving our deep learning-based classification.
To deploy Cuckoo Sandbox effectively, specific system requirements must be met. The host operating system should be a Linux distribution (e.g., Ubuntu 18.04 LTS, Debian, or CentOS) due to its stability and compatibility with virtualization tools. Hardware demands include a multi-core CPU (e.g., Intel i7 or equivalent) to support concurrent VM execution, ample RAM (minimum 8 GB, recommended 16 GB or more) to accommodate multiple guests, and substantial disk space (at least 500 GB, ideally SSD-backed) to store malware samples, VM images, and analysis logs. Virtualization software such as VirtualBox, VMware Workstation, or KVM is mandatory to create the isolated execution environments, with configuration options to emulate diverse Windows versions (e.g., Windows 10, 7) matching target malware’s operational context. Additionally, Python 3.x is required, along with dependencies like libvirt and volatility, to support Cuckoo’s scripting and memory forensics capabilities. Proper network configuration—including a virtual NIC bridged to a monitored subnet—ensures safe capture of network-based behaviors without risking external exposure.
Cuckoo Sandbox’s strengths lie in its modularity and extensibility. It supports plugins for custom analysis (e.g., memory dump parsing with Volatility), integrates with external services like VirusTotal for signature verification, and scales to handle large sample volumes via distributed setups. However, its reliance on virtualization introduces potential weaknesses, such as detection by VM-aware malware that alters behavior when sandboxed (e.g., checking for hypervisor artifacts like CPUID instructions). To mitigate this, advanced configurations employ bare-metal sandboxes or anti-evasion techniques (e.g., randomizing system artifacts), though these fall beyond our current scope.
In our research, Cuckoo Sandbox serves as the linchpin for dynamic PE malware analysis, enabling the extraction of runtime behaviors critical to our deep learning and genetic algorithm framework. By isolating and dissecting malware in this controlled environment, we obtain the raw data—API call sequences—that fuel our $n$-gram feature extraction, classification, and concept drift handling, as elaborated in subsequent sections.